# Automorphisms of relatively hyperbolic groups and the Farrell–Jones Conjecture

### 23 Nov 2023 v1

@@ 1. Introduction
is an isomorphism. Here $\underline{\underline{E}}G$ is the classifying space for the family of virtually cyclic subgroups, $\mathbf{K}_{R}$ is the algebraic $K$-theory spectrum for the ring $R$, and $K_{n}(RG)$ is the algebraic $K$-theory of the group ring $RG$. There are variants of the conjecture for Waldhausen’s $A$-theory and for $L$-theory. The conjecture for $L$-theory, as well as a detailed account of the Farrell–Jones Conjecture, and the objects involved can be found in W. Lück’s book project [Lüc]. For recent progress on $A$-theory the reader should consult [${}^{+}$18].
@@
Computing the algebraic $K$-theory of a group ring $RG$ is a very difficult problem. In principle, knowing that FJC holds for $G$ gives a method of computing $K_{n}(RG)$ using equivariant algebraic topology. It also has a number of other applications, for example, to the Borel Conjecture [BL12] and to computing the Whitehead group $\mathrm{Wh}(G)$. Knowledge of $\mathrm{Wh}(G)$ is a fundamental step in classifications of higher dimensional manifolds with fundamental group $G$.
@@
From this point onward, by the Farrell–Jones Conjecture for $G$, we mean the most general setting, that is, the fibred Farrell–Jones Conjecture with respect to the family of virtually cyclic subgroups $\mathcal{VC}$. See for example [${}^{+}$18] and [Lüc] for a discussion of these terms. We will denote the classes of group satisfying the FJC for $X$-theory by $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$ where $X$ is $A$-, $K$-, or $L$-theory.
@@
The class $\operatorname{\mathbf{FJC}}_{\mathbf{K}}$ of groups known to satisfy FJC for algebraic $K$-theory is large: containing hyperbolic groups [BLR08a], many relatively hyperbolic groups [Bar17], $\mathrm{CAT}(0)$ groups [Weg12] (see also [BL12] and [KR17]), soluble groups [Weg15], $\mathrm{GL}_{n}(\mathbb{Z})$ [BLRR14] and more generally lattices in connected Lie groups [BFL14] and $S$-arithmetic groups [Rüp16], as well as mapping class groups [BB19], normally poly-free groups [BKW21a], and suspensions of virtually torsion free hyperbolic groups [BFW23]. The class enjoys many closure properties: it passes to arbitrary subgroups, finite index overgroups, and directed colimits. For more information the reader is referred to the surveys [BLR08b, LR05, Lüc10, Bar16].
@@
One property that is not known is whether $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$ is closed under extensions $1\to N\to\Gamma\to Q\to 1$. One direction of interest is to put conditions on $N$ so that $\Gamma$ is in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$ whenever $Q$ is. By [BFL14, Theorem 2.7] and [${}^{+}$18, Theorem 1.1(ii)] this reduces to understanding cyclic extensions of $G$; which is to say the suspensions $N_{\Phi}=N\rtimes_{\Phi}\mathbb{Z}$, where $\Phi$ is some automorphism of $N$ defining this suspension.
@@
Intuitively, a group $G$ is hyperbolic relative to ${\mathcal{P}}$ if its geometry is hyperbolic “away from the subgroups $P\in{\mathcal{P}}$.” One (of many: see [Hru10] for the definitions as well as proofs of their equivalence) way to formalise this uses the notion of coning off a Cayley graph: take a vertex for every coset $gP$ of each element of $P$, and add an edge from each element of $gP$ to the new vertex. The group $G$ is hyperbolic relative to ${\mathcal{P}}$ if the resulting graph is $\delta$-hyperbolic in the sense of Gromov, and fine: every edge is contained in finitely many cycles of a given finite length. An automorphism of $G$ lies in the subgroup $\operatorname{\mathrm{Aut}}(G,{\mathcal{P}})$ if it preserves the conjugacy classes of every subgroup $P\in{\mathcal{P}}$. For more information on $\operatorname{\mathrm{Aut}}(G,{\mathcal{P}})$ see [MO12] and [GL15].
@@
Recently, Bestvina, Fujiwara and Wigglesworth [BFW23] proved the suspension of a virtually torsion free hyperbolic group satisfies the Farrell–Jones conjecture. We extend this result to a large class of relatively hyperbolic groups.
A.
@@
Let $(G,{\mathcal{P}})$ be a virtually torsion-free or one-ended relatively hyperbolic group with ${\mathcal{P}}$ finite and let $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{P}})$. If for every $[P]\in{\mathcal{P}}$ we have $P_{\Phi}\in\operatorname{\mathbf{FJC}_{\mathbf{X}}}$, then $G_{\Phi}\in\operatorname{\mathbf{FJC}_{\mathbf{X}}}$.
@@
These hypotheses include, for instance, all suspensions of toral relatively hyperbolic groups and more generally one-ended or virtually torsion-free groups that are hyperbolic relative to virtually polycyclic or soluble subgroups. Note that this removes the assumption of virtual torsion-freeness in [BFW23] for one-ended hyperbolic groups. This is pertinent since it is a well known question of Gromov whether every hyperbolic group is residually finite (and hence virtually torsion-free).
@@
With infinitely ended groups more care is needed, we discuss this further in Section 1.B.
@@ 1.A. Applications
The assumption that peripheral subgroups are NRH is needed for B, since it requires A to hold for arbitrary automorphisms. The key point being that $\operatorname{\mathrm{Aut}}(N;{\mathcal{P}})$ has finite index in $\operatorname{\mathrm{Aut}}(N)$ under this extra hypothesis.
It is a major open problem whether $\operatorname{\mathrm{Out}}(F_{N})$ satisfies $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$. Whilst we do not solve this, using A we are able to show automorphism groups of one-ended groups hyperbolic relative to virtually polycyclic groups satisfy FJC. In particular, $\operatorname{\mathrm{Aut}}(G)$ and $\operatorname{\mathrm{Out}}(G)$ for $G$ a one-ended hyperbolic group satisfy $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$.
C.
If $G$ is a one ended group hyperbolic relative to finitely many conjugacy classes of virtually polycyclic groups, then $\operatorname{\mathrm{Aut}}(G)$ and $\operatorname{\mathrm{Out}}(G)$ are in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$.
@@ 1.B. Remarks on the proofs
As is usual for (relatively) hyperbolic groups, there are two main flavours to our arguments, depending on the number of ends of $G$. In both cases we apply a result of Knopf [Kno19] allowing us to deduce that a group acting acylindrically on a tree satisfies the Farrell–Jones conjecture if and only if its vertex groups do, though the source of the trees is different in each case.
@@
For one-ended relatively hyperbolic groups, we have access to the powerful machinery of JSJ decompositions developed (in this generality) by Guirardel and Levitt [GL17]. We consider three related trees: the canonical JSJ decomposition $T^{\mathrm{can}}$ relative to the peripheral subgroups ${\mathcal{P}}$, a refinement $T^{\phi}$ of $T^{\mathrm{can}}$ which better suits the study of an outer automorphism $\phi$ and another tree that we call $T^{\mathrm{Per}}$. This tree is the canonical JSJ tree relative to the (non-elementary) periodic subgroups of the outer automorphism $\phi$. That is, we require that the periodic subgroups of every representative $\Phi\operatorname{ad}_{g}$, are elliptic. Our main structural result about this tree is 5.19: even without assuming that the periodic subgroups are finitely generated, they agree exactly with the rigid vertices of $T^{\mathrm{Per}}$. We prove that this ensures that the induced action of $G\rtimes_{\Phi}\mathbb{Z}$ is acylindrical, and then analyse vertex groups that can appear in this new action.
@@
We remark that the strong uniqueness properties of the JSJ decomposition imply that for a one-ended, torsion free hyperbolic group, the rigid vertex groups of the tree considered in [BFW23] agree with those in our $T^{\mathrm{Per}}$.
@@
We consider the case when $G$ is infinitely ended and has a finite index subgroup which is a free product of one-ended groups. Being virtually torsion free is sufficient but not necessary for this to occur, and we provide necessary and sufficient conditions in 8.1. Once we have a free product splitting we have the following combination-type theorem.
D.
@@
Let $G=G_{1}\ast\ldots\ast G_{k}\ast F_{N}$ be a free product of finitely generated groups, let =${\mathcal{F}}^{\prime}=\{[G_{1}],\ldots,[G_{k}]\}$ and let $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{F}}^{\prime})$. For every $i\in\{1,\ldots,k\}$, denote by $\Phi_{i}$ an element of the outer class of $\Phi$ preserving $G_{i}$. If for every $i\in\{1,\ldots,k\}$ the group $G_{i}\rtimes_{\Phi_{i}}\mathbb{Z}$ is in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$, then $G\rtimes_{\Phi}\mathbb{Z}$ is in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$.
@@
D is proved by induction on the Grushko rank $k+N$. There are two kinds of induction step, depending on whether the maximal periodic free factor system is sporadic or not. A free factor system $(G,{\mathcal{F}})$ is sporadic if $G\cong G_{1}\ast G_{2}$ or $G\cong G_{1}\ast\mathbb{Z}$. This division might seem unusual to experts; a more standard division (for instance, in [BFW23] as well as throughout the $\operatorname{\mathrm{Out}}(F_{n})$ literature) depends instead on if the automorphism is polynomially or exponentially growing. Polynomially growing automorphisms are always sporadic in this sense, but so are some exponentially growing automorphisms. The non-sporadic case uses Dahmani and Li’s work on relative hyperbolicity for suspensions of free factors [DL22], whereas in the sporadic case we use the fact that these splittings are rigid.
@@
These rigidity arguments hold equally well for sporadic Stallings–Dunwoody decompositions, and so we are still able to obtain some results without first passing to a finite index free product: see 8.3.
@@
Some previous results of this flavour have concluded the stronger property that the suspension is in the class $\operatorname{\mathcal{AC}}(\operatorname{\mathbf{VNil}})$. Every group in this class satisfies the Farrell–Jones conjecture [BB19]. This class has similar closure properties to the class of groups satisfying the Farrell–Jones conjecture, except that $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$ is closed under directed colimits while $\operatorname{\mathcal{AC}}(\operatorname{\mathbf{VNil}})$ is not known to be. However, for the majority of the paper we work directly with the class $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$.
@@
E.
Suppose $(G,{\mathcal{P}})$ is one-ended or virtually torsion free, and hyperbolic relative to finitely many conjugacy classes of slender subgroups. Then for every automorphism $\Phi$ of $G$, $\Gamma\coloneqq G\rtimes_{\Phi}\mathbb{Z}$ is in $\operatorname{\mathcal{AC}}(\operatorname{\mathbf{VNil}})$.
Except for replacing each periodic subgroup with the fixed subgroup of a power, the proof of this theorem is identical to the proof of A. We discuss this in a little more detail after completing that proof.
@@
Following work of Bunke, Kaprowski, and Winges [BKW21b] our results apply equally well to the Farrell–Jones Conjecture for localising invariants, that is, with coefficients in $H\colon\mathbf{Cat}_{\infty,\ast}^{\mathrm{Lex}}\to\mathbf{M}$ a lax monoidal finitary localising invariant with values in a stably monoidal and cocomplete stable $\infty$-category which admits countable products. We refer the reader to the introduction of loc. cit. for more information.
@@ 1.D. Structure of the paper
Section 2 introduces the relevant background results on the Farrell–Jones conjecture.
Section 3 contains definitions and results on free products and their automorphisms, needed for Section 6.
Section 4 collects results on JSJ decompositions of one ended relatively hyperbolic groups, and provides a lemma on acylindricity when passing to the action of a suspension.
The one-ended case of A is proved in Section 5 by careful analysis of a certain JSJ tree. From this analysis, we also deduce 5.22.
In Section 6 we prove D and the infinitely-ended case of A. Using these results we prove A and E.
In Section 7, we deduce C from A.
Finally, in Section 8 we extend A as far as possible with our current techniques to groups which are infinitely ended but do not split as free products. The arguments of the last three sections are almost independent of Sections 4 and 5, apart from requiring the background information on trees of cylinders from Section 4.A.

### 02 Mar 2026 v2

# arXiv:2311.14036v2[math.KT] 02 Mar 2026

@@ 1. Introduction
is an isomorphism. Here $\underline{\underline{E}}G$ is the classifying space for the family of virtually cyclic subgroups, $\mathbf{K}_{R}$ is the algebraic $K$-theory spectrum for the ring $R$, and $K_{n}(RG)$ is the algebraic $K$-theory of the group ring $RG$. There are variants of the conjecture for Waldhausen’s $A$-theory and for $L$-theory. The conjecture for $L$-theory, as well as a detailed account of the Farrell–Jones Conjecture, and the objects involved can be found in W. Lück’s book project [41]. For recent progress on $A$-theory the reader should consult [18].
@@
Computing the algebraic $K$-theory of a group ring $RG$ is a very difficult problem. In principle, knowing that FJC holds for $G$ gives a method of computing $K_{n}(RG)$ using equivariant algebraic topology. It also has a number of other applications, for example, to the Borel Conjecture [6] and to computing the Whitehead group $\mathrm{Wh}(G)$. Knowledge of $\mathrm{Wh}(G)$ is a fundamental step in classifications of higher dimensional manifolds with fundamental group $G$.
@@
From this point onward, by the Farrell–Jones Conjecture for $G$, we mean the most general setting, that is, the fibred Farrell–Jones Conjecture with respect to the family of virtually cyclic subgroups $\mathcal{VC}$. See for example [18] and [41] for a discussion of these terms. We will denote the classes of groups satisfying the FJC for $X$-theory by $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$ where $X$ is $A$-, $K$-, or $L$-theory.
@@
The class $\operatorname{\mathbf{FJC}}_{\mathbf{K}}$ of groups satisfying FJC for algebraic $K$-theory is large: containing hyperbolic groups [5], many relatively hyperbolic groups [8], $\mathrm{CAT}(0)$ groups [50] (see also [6] and [35]), soluble groups [51], $\mathrm{GL}_{n}(\mathbb{Z})$ [3] and more generally lattices in connected Lie groups [1] and $S$-arithmetic groups [48], as well as mapping class groups [2], normally poly-free groups [14], and suspensions of virtually torsion free hyperbolic groups [11]. The class enjoys many closure properties: it passes to arbitrary subgroups, finite index overgroups, and directed colimits. For more information the reader is referred to the surveys [4, 40, 42, 7].
@@
One property that is not known is whether $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$ is closed under extensions $1\to N\to\Gamma\to Q\to 1$. One direction of interest is to put conditions on $N$ so that $\Gamma$ is in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$ whenever $Q$ is. By [1, Theorem 2.7] and [18, Theorem 1.1(ii)] this reduces to understanding cyclic extensions of $G$; which is to say the suspensions $N_{\Phi}=N\rtimes_{\Phi}\mathbb{Z}$, where $\Phi$ is some automorphism of $N$ defining this suspension.
@@
Intuitively, a group $G$ is hyperbolic relative to $\mathcal{P}$} if its geometry is hyperbolic “away from the subgroups $P\in{\mathcal{P}}$.” One (of many: see [33] for the definitions as well as proofs of their equivalence) way to formalise this uses the notion of coning off a Cayley graph: take a vertex for every coset $gP$ of each element of $P$, and add an edge from each element of $gP$ to the new vertex. The group $G$ is hyperbolic relative to $\mathcal{P}$} if the resulting graph is $\delta$-hyperbolic in the sense of Gromov, and fine: every edge is contained in finitely many cycles of a given finite length. An automorphism of $G$ lies in the subgroup $\operatorname{\mathrm{Aut}}(G,{\mathcal{P}})$ if it preserves the conjugacy classes of every subgroup $P\in{\mathcal{P}}$. For more information on $\operatorname{\mathrm{Aut}}(G,{\mathcal{P}})$ see [44] and [27].
@@
Recently, Bestvina, Fujiwara and Wigglesworth [11] proved the suspension of a virtually torsion free hyperbolic group satisfies the Farrell–Jones conjecture. We extend this result to a large class of relatively hyperbolic groups.
˜A.
@@
Let $(G,{\mathcal{P}})$ be a virtually torsion-free or one-ended relatively hyperbolic group with $\mathcal{P}$} finite and let $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{P}})$. If for every $P]\in{\mathcal{P}}$ we have $P_{\Phi}\in\operatorname{\mathbf{FJC}_{\mathbf{X}}}$, then $G_{\Phi}\in\operatorname{\mathbf{FJC}_{\mathbf{X}}}$.
@@
These hypotheses include, for instance, all suspensions of toral relatively hyperbolic groups and more generally one-ended or virtually torsion-free groups that are hyperbolic relative to virtually polycyclic or soluble subgroups. Note that this removes the assumption of virtual torsion-freeness in [11] for one-ended hyperbolic groups. This is pertinent since it is a well known question of Gromov whether every hyperbolic group is residually finite (and hence virtually torsion-free).
@@
With infinitely ended groups more care is needed, we discuss this further in Section˜1.B.
@@ 1.A. Applications
The assumption that peripheral subgroups are NRH is needed for ˜B, since it requires ˜A to hold for arbitrary automorphisms. The key point being that $\operatorname{\mathrm{Aut}}(N;{\mathcal{P}})$ has finite index in $\operatorname{\mathrm{Aut}}(N)$ under this extra hypothesis.
It is a major open problem whether $\operatorname{\mathrm{Out}}(F_{N})$ satisfies $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$. Whilst we do not solve this, using ˜A we are able to show automorphism groups of one-ended groups hyperbolic relative to virtually polycyclic groups satisfy $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$. In particular, $\operatorname{\mathrm{Aut}}(G)$ and $\operatorname{\mathrm{Out}}(G)$ for $G$ a one-ended hyperbolic group satisfy $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$.
˜C.
@@ 1.B. Remarks on the proofs
As is usual for (relatively) hyperbolic groups, there are two main flavours to our arguments, depending on the number of ends of $G$. In both cases we apply a result of Knopf [37] allowing us to deduce that a group acting acylindrically on a tree satisfies the Farrell–Jones conjecture if and only if its vertex groups do, though the source of the trees is different in each case.
@@
For one-ended relatively hyperbolic groups, we have access to the powerful machinery of JSJ decompositions developed (in this generality) by Guirardel and Levitt [30]. We consider three related trees: the canonical JSJ decomposition $T^{\mathrm{can}}$ relative to the peripheral subgroups $\mathcal{P}$}, a refinement $T^{\phi}$ of $T^{\mathrm{can}}$ which better suits the study of an outer automorphism $\phi$ and another tree that we call $T^{\mathrm{Per}}$. This tree is the canonical JSJ tree relative to the (non-elementary) periodic subgroups of the outer automorphism $\phi$. That is, we require that the periodic subgroups of every representative $\Phi\operatorname{ad}_{g}$, are elliptic. Our main structural result about this tree is ˜5.21: even without assuming that the periodic subgroups are finitely generated, they agree exactly with the rigid vertices of $T^{\mathrm{Per}}$. We prove that this ensures that the induced action of $G\rtimes_{\Phi}\mathbb{Z}$ is acylindrical, and then analyse vertex groups that can appear in this new action.
@@
We remark that the strong uniqueness properties of the JSJ decomposition imply that for a one-ended, torsion free hyperbolic group, the rigid vertex groups of the tree considered in [11] agree with those in our $T^{\mathrm{Per}}$.
@@
We consider the case when $G$ is infinitely ended and has a finite index subgroup which is a free product of one-ended groups. Being virtually torsion free is sufficient but not necessary for this to occur. Once we have a free product splitting we have the following combination-type theorem.
˜D.
@@
Let $G=G_{1}\ast\ldots\ast G_{k}\ast F_{N}$ be a free product of finitely generated groups, let ${\mathcal{F}}^{\prime}=\{[G_{1}],\ldots,[G_{k}]\}$ and let $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{F}}^{\prime})$. For every $i\in\{1,\ldots,k\}$, denote by $\Phi_{i}$ an element of the outer class of $\Phi$ preserving $G_{i}$. If for every $i\in\{1,\ldots,k\}$ the group $G_{i}\rtimes_{\Phi_{i}}\mathbb{Z}$ is in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$, then $G\rtimes_{\Phi}\mathbb{Z}$ is in $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$.
@@
˜D is proved by induction on the Grushko rank $k+N$. There are two kinds of induction step, depending on whether the maximal periodic free factor system is sporadic or not. A free factor system $(G,{\mathcal{F}})$ is sporadic if $G\cong G_{1}\ast G_{2}$ or $G\cong G_{1}\ast\mathbb{Z}$. This division might seem unusual to experts; a more standard division (for instance, in [11] as well as throughout the $\operatorname{\mathrm{Out}}(F_{n})$ literature) depends instead on if the automorphism is polynomially or exponentially growing. Polynomially growing automorphisms are always sporadic in this sense, but so are some exponentially growing automorphisms. The non-sporadic case uses Dahmani and Li’s work on relative hyperbolicity for suspensions of free factors [16], whereas in the sporadic case we use the fact that these splittings are rigid.
@@
These rigidity arguments hold equally well for sporadic Stallings–Dunwoody decompositions, and so we are still able to obtain some results without first passing to a finite index free product: see ˜8.1.
@@
Some previous results of this flavour have concluded the stronger property that the suspension is in the class $\operatorname{\mathcal{AC}}(\operatorname{\mathbf{VNil}})$. This class consists of groups having particularly nice actions on compact Euclidean retracts, see [2] for a precise definition. Every group in this class satisfies the Farrell–Jones conjecture [2]. This class has similar closure properties to the class of groups satisfying the Farrell–Jones conjecture, except that $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$ is closed under directed colimits while $\operatorname{\mathcal{AC}}(\operatorname{\mathbf{VNil}})$ is not known to be. However, for the majority of the paper we work directly with the class $\operatorname{\mathbf{FJC}_{\mathbf{X}}}$.
@@
˜E.
Suppose $(G,{\mathcal{P}})$ is one-ended or virtually torsion free, and hyperbolic relative to finitely many conjugacy classes of virtually polycyclic subgroups. Then for every automorphism $\Phi$ of $G$, $\Gamma\coloneqq G\rtimes_{\Phi}\mathbb{Z}$ is in $\operatorname{\mathcal{AC}}(\operatorname{\mathbf{VNil}})$.
Except for replacing each periodic subgroup with the fixed subgroup of a power, the proof of this theorem is identical to the proof of ˜A. We discuss this in a little more detail after completing that proof.
@@
Following work of Bunke, Kaprowski, and Winges [15] our results apply equally well to the Farrell–Jones Conjecture for localising invariants, that is, with coefficients in $H\colon\mathbf{Cat}_{\infty,\ast}^{\mathrm{Lex}}\to\mathbf{M}$ a lax monoidal finitary localising invariant with values in a stably monoidal and cocomplete stable $\infty$-category which admits countable products. We refer the reader to the introduction of loc. cit. for more information.
@@ 1.D. Structure of the paper
Section˜2 introduces the relevant background results on the Farrell–Jones conjecture.
Section˜3 contains definitions and results on free products and their automorphisms, needed for Section˜6.
Section˜4 collects results on JSJ decompositions of one-ended relatively hyperbolic groups, and provides a lemma on acylindricity when passing to the action of a suspension.
The one-ended case of ˜A is proved in Section˜5 by careful analysis of a certain JSJ tree. From this analysis, we also deduce ˜5.24.
In Section˜6 we prove ˜D and the infinitely-ended case of ˜A. Using these results we prove ˜A and ˜E.
In Section˜7, we deduce ˜C from ˜A.
Finally, in Section˜8 we extend ˜A as far as possible with our current techniques to groups which are infinitely ended but do not split as free products. The arguments of the last three sections are almost independent of Sections˜4 and 5, apart from requiring the background information on trees of cylinders from Section˜4.A.